OpenStudy (anonymous):

An inequality is shown below: -np - 5 ≤ 4(c - 2) Which of the following solves for n? @amistre64 @ganeshie8 @satellite73 @SolomonZelman

OpenStudy (amistre64):

do we assume n is postive to start with?

OpenStudy (anonymous):

n ≥ - the quantity 4 times c minus 3 all over p n ≥ - the quantity 4 times c minus 13 all over p n ≤ - the quantity 4 times c minus 3 all over p n ≤ - the quantity 4 times c minus 13 all over p

OpenStudy (amistre64):

if we have -a < b, and want to solve for a, its best to add it over 0 < b + a, therefore: -b < a this avoids alot of headache trying to remember to flip signs

OpenStudy (anonymous):

@amistre64 but my problem is much more complex

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

yes

OpenStudy (perl):

-np - 5<=4(c - 2) -np - 5 <= 4c - 8 -np <= 4c -8 + 5 -np <= 4c - 3 did you get up to here?

OpenStudy (perl):

ok if you divide by a negative, you have to change the sign of the inequality

OpenStudy (anonymous):

so n<=4c-3/p

OpenStudy (perl):

-np <= 4c-3 n >= - (4c - 3 ) / p

OpenStudy (perl):

you just need the minus in front

OpenStudy (anonymous):

so the answer is A

OpenStudy (perl):

sorry i can't give the answer :)

OpenStudy (amistre64):

is p assumed to be positive to start with?

OpenStudy (amistre64):

lets see if it matter: -n(-3) > 20 n < 20/-(-3) n < 20/3 < 6 1/3 so, let n=-1 since -1 < 6 -(-1)(-3) > 20 is this true or false?

OpenStudy (anonymous):

true

OpenStudy (amistre64):

you think -3 is bigger than 20?

OpenStudy (anonymous):

oh wait no i had my sign messed up

OpenStudy (amistre64):

when we divide by a variable, we have to have some sort of restrictions on it in an inequality; since a negative value flips the equailty sign

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the answer is eaither a or b because thats the ones that the sign flips

OpenStudy (amistre64):

-np - 5 <= 4(c - 2) -4(c-2) - 5 <= np ------------------- if p is positive then [-4(c-2) - 5]/p <= n ------------------- if p is negative then [-4(c-2) - 5]/p >= n

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

i spose its generally assumed the p is positive tho

OpenStudy (anonymous):

so its A